# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@ B.3 Proof of Proposition 1
We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ would lead to model that is not observationally equivalent with the original one. Specifically, we say that two models are observational equivalent if and only if their joint densities are equal. After that, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified.
@@
Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$. It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is an one-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{2}$ in the region where the gradient of $f(x_{1},x_{2})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $f(x_{1},x_{2})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{2}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{2}$. That is, $st(x_{1};\nu_{i},\lambda_{i})=st(x_{2};\nu_{j},\lambda_{j})$ only on sets of zero Lebesgue measure, concluding that $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$.
@@
Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model. Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{T},...,y_{0})$. Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
@@
where $\mathcal{T}_{m}$, $m=1,...,M$, is the (nonempty) set of time indices (identified by Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
@@
By Corollary B.1, different orderings and signs of the elements of $B_{m}^{-1}(y_{t}-\mu_{y,t})$ yield different values of $\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ under Conditions (b) and (c) of Proposition 1. Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$. Therefore, as the shocks and thereby $y_{t}$ follows a continuous distribution, we have
$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.19)
@@
almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$. Since this result holds almost everywhere in $Y_{T}\in\mathbb{R}^{T}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.
By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$. Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns. Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed. Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different values of the joint density function of the model almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$, i.e., to models that are not observationally equivalent. Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $Y_{T}\in\mathbb{R}^{dT}$.∎
@@ B.4 Proof of Proposition 2
To proof consists of two parts. First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix. In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$. Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified. For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$, and $z_{t}\in\mathbb{R}^{T}$ (where $z_{t}$ is the switching variable), whereas for models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.
By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$. With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given. In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
@@
At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{m,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.21)
@@
By Condition (B), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, Equation (B.23) is nontrivial, as by invertibility, none of the columns of $B_{m}$ can be equal or negatives of each other. Hence, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.23) forms a subspace in $\mathbb{R}^{d\times d(M-1)}$ with the dimension of this subspace being less than $(M-1)d^{2}$, as it must satisfy a nontrivial linear equation. Consequently, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.23) has a Lebesgue measure zero in $[B_{2}:\cdots:B_{M}]\in\mathbb{R}^{d\times d(M-1)}$, implying that the set of matrices $B_{1},...,B_{M}$ satisfying this equation have a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
Suppose then $D_{t}P_{t}\neq I_{d}$ and denote the left side of Equation (B.22) as $C_{t}\in\mathbb{R}^{d\times d}$. By Condition (B), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, the right side of Equation (B.22) is nonzero, thus, implying that also $C_{t}$ is not a matrix zeros. Equation (B.22) is thereby nontrivial and can be expressed as
@@
By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.22) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.23) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Consider the impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ with some invertible $(d\times d)$ matrices $B_{1},...,B_{M}$. By Lemma 2, $B_{y,t}$ is uniquely identified at each $t$ up to ordering and signs of its columns (conditionally on $\mathcal{F}_{t-1}$). Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A)), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$. Hence, reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ would lead to a model that is not observationally equivalent with the original one. The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A), implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$). Thus, $B_{y,t}$ is uniquely identified at all $t$.
@@
Denote the $ij$th element of $\tilde{B}_{m}$ as $\tilde{b}_{ij}^{(m)}$. At each $t$, Equation (B.25) then defines a system of linear equations involving $d^{2}$ equations and $Md^{2}$ variables $\tilde{b}_{ij}^{(m)}$, $j,i=1,...,d$, $m=1,...,M$. The weights $\alpha_{m,t}$ are treated as constant multipliers here, as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here. If $\alpha_{m,t}=0$ for some regimes $m$, the corresponding variables in $\tilde{B}_{m}$ drop out of the system of equations, however.
Under Condition (C), the vector $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors $(\alpha_{n_{1},t},...,\alpha_{n_{M^{\circ}},t})$, $n_{l}\in\mathbb{M^{\circ}}$, $\mathbb{M^{\circ}}\subset\{1,...,M\}$, where $M^{\circ}\geq 2$ denotes the number of elements in $\mathbb{M^{\circ}}$, take at least $M+1$ linearly independent values (plus one accounting for transition weights that equal to zero). Therefore, Equation (B.25) defined together for all $t$ defines a system of linear equations involving $Md^{2}$ variables and at least $Md^{2}$ linearly independent equations. Moreover, as $(\alpha_{1,t},...,\alpha_{M,t})$ and all its subvectors take at least $M+1$ linearly independent values, each variable $\tilde{b}_{ij}^{(m)}$ appears in at least $M$ linearly independent the equations. Thus, there exists at most one solution to this system of linear equations. Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywehere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified. As already stated, the transition weights weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$ for logistic models and almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ for models with nonrandom exogenous weight, concluding the proof.∎

### v4

@@ B.3 Proof of Proposition 1
We start by showing that under the conditions of Proposition 1, the ordering and signs of the columns of $B_{y,t}$ are fixed for all $t$ in the sense that models before and after reordering or changing the signs of the columns of $B_{y,t}$ at any $t$ are not observationally equivalent. Specifically, we say that two models are observational equivalent if and only if their joint densities are equal. After that, we combine this result with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified.
@@
Denote $f(x_{1},x_{2})\equiv st(x_{1};\nu_{i},\lambda_{i})-st(x_{2};\nu_{j},\lambda_{j})$, which is not identically zero by $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$. It is straightforward to show (by induction) that the function $st(x_{i};\nu_{i},\lambda_{i})$, and also thereby $f(x_{1},x_{2})$, is infinitely continuously differentiable. It then follows from Implicit function theorem that in the points $x^{0}=(x_{1,0},x_{2,0})\in\mathbb{R}^{2}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, the set $\mathcal{X}\equiv\{(x_{1},x_{2})\in\mathbb{R}^{2}|f(x_{1},x_{2})=0\}$ is a one-dimensional manifold locally in a neighbourhood of $x^{0}$. Because this argument applies at all points $x^{0}$ where the gradient of $f(x_{1},x_{2})$ is nonzero, it follows that the set $\mathcal{X}$ is a one-dimensional manifold in regions where this condition holds. Therefore, Lebesque measure of the set $\mathcal{X}$ is zero in $\mathbb{R}^{2}$ in the region where the gradient of $f(x_{1},x_{2})$ is nonzero. On the other hand, by Sard’s theorem (Sard, 1994), the set of points where the gradient of $f(x_{1},x_{2})$ is zero has a Lebesgue measure zero in $\mathbb{R}^{2}$. Since the union of two sets with a Lebesgue measure zero has a Lebesgue measure zero, these results can be combined to conclude that the $\mathcal{X}$ has a Lebesgue measure zero in $\mathbb{R}^{2}$. That is, $st(x_{1};\nu_{i},\lambda_{i})=st(x_{2};\nu_{j},\lambda_{j})$ only on sets of zero Lebesgue measure, concluding that $st(x_{1};\nu_{i},\lambda_{i})\neq st(x_{2};\nu_{j},\lambda_{j})$ almost everywhere in $(x_{1},x_{2})\in\mathbb{R}^{2}$.
@@
Condition (a) of Proposition 1 fixes the ordering and signs of the columns of $B_{1}$, so it suffices to show that under Conditions (b) and (c), changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ would change the joint density of the model. Suppose the observed time series is $y_{t-p+1},...,y_{0},y_{1},...,y_{T}$ and denote $Y_{T}=(y_{1},...,y_{T})$. Conditional on the initial values $y_{t-p+1},...,y_{0}$, the joint density of the model is
@@
where $\mathcal{T}_{m}$, $m=1,...,M$, is the set of time indices (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and nonempty by Condition (3) of Lemma 2) that satisfy $r_{m-1}<z_{t}\leq r_{m}$ for the regime $m$, $\mu_{y,t}$ is the conditional mean of the model (identified by Lemma 2 almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$), and the dependence of the density functions $f_{Y}(\cdot;B_{1},...,B_{M})$ and $f_{y}(\cdot;B_{m})$ on the impact matrices $B_{1},...,B_{M}$ is now made explicit in the notation (but their dependence on the other parameters is omitted).
@@
By Corollary B.1, different orderings and signs of the elements of $B_{m}^{-1}(y_{t}-\mu_{y,t})$ yield different values of $\prod_{i=1}^{d}st(I_{d,i}^{\prime}B_{m}^{-1}(y_{t}-\mu_{y,t});\nu_{i},\lambda_{i})$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ under Conditions (b) and (c) of Proposition 1. Moreover, as $|\det(B_{m})|$ is invariant to the ordering and signs of the columns of $B_{m}$, it follows from that, $f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m}))\neq f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))$ almost everywhere in $y_{t}\in\mathbb{R}^{d}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$. Therefore, as $\mathcal{T}_{1},...,\mathcal{T}_{M}$ and $\mu_{y,t}$, $t=1,...,T$ are identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ by Lemma 2 and the shocks and thereby $y_{t}$ follow continuous distributions, we have
$\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\boldsymbol{\tau}_{m}))\neq\prod_{t\in\mathcal{T}_{m}}f_{y}(y_{t}-\mu_{y,t};B_{m}(\tilde{\boldsymbol{\tau}}_{m})),$ (B.23)
@@
almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$ whenever $\tilde{\boldsymbol{\tau}}_{m}\neq\boldsymbol{\tau}_{m}$ or $\tilde{\boldsymbol{\iota}}_{m}\neq\boldsymbol{\iota}_{m}$ for any $m=2,...,M$. Since this result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, and $B_{y,t}$ is here invertible for all invertible $B_{1},...,B_{M}$, the result holds almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.
By Lemma 2, $B_{y,t}$ is identified up to ordering and signs of its columns at each $t$. Thus, each impact matrix $B_{1},...,B_{M}$ of the regimes is identified up to ordering signs of its columns. Under the conditions of Proposition 1, the ordering and signs of the columns of $B_{1}$ are directly fixed. Moreover, by the above discussion, different permutations or sign changes of the columns of any of $B_{2},...,B_{M}$ lead to different joint densities of the model almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$, i.e., to models that are not observationally equivalent. Thus, only models corresponding to the same permutations and sign changes of the columns of $B_{2},...,B_{M}$ are observationally equivalent, implying that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},Y_{T})\in\mathbb{R}^{d(T+p)}$.∎
@@ B.4 Proof of Proposition 2
The proof consists of two parts. First, we show that if the ordering and signs of the columns of $B_{1}$ is fixed, changing the ordering or signs of the columns of any of $B_{2},...,B_{M}$ leads to an impact matrix $B_{y,t}$ that cannot, at all $t$, be obtained by reordering or changing the signs of the columns of the original impact matrix. In other words, we show that fixing the ordering signs of the columns of $B_{1}$ effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ and thereby of $B_{y,t}$. Then, we make use of this result together with the result of Lemma 2 (that $B_{y,t}$ is identified up ordering and signs of its columns at each $t$) to show that the matrices $B_{1},...,B_{M}$ are uniquely identified. For models with logistic transition weights (2.4), the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. For models with nonrandom exogenous transition weights, the result holds almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$.
By Lemma 2, for logistic models (2.4) (with $M=2$), the parameters $c,\gamma$, and threreby the transition weights $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},$ $y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$). With nonrandom exogenous transition weights, the identification of $\alpha_{1,t},...,\alpha_{M,t}$$t=1,...,T$ is given. In either case, the transition weights can be treated as uniquely identified in the rest of the proof, where, for readability, we do not repeatedly remind that the identification holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$ for exogenous $z_{t}$).
@@
Suppose $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$. At each time period $t$, the impact matrix $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ is obtained by reordering or changing the signs of the columns of $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ if and only if
$\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\boldsymbol{\tau}_{m},\boldsymbol{\iota}_{m})=(\alpha_{1,t}B_{1}(\boldsymbol{\tau}_{1},\boldsymbol{\iota}_{1})+\sum_{m=2}^{M}\alpha_{m,t}B_{m}(\tilde{\boldsymbol{\tau}}_{m},\tilde{\boldsymbol{\iota}}_{m}))D_{t}P_{t}$ (B.25)
@@
By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, Equation (B.27) is nontrivial, as by invertibility, none of the columns of $B_{m}$ can be equal or negatives of each other. Hence, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.27) forms a subspace in $\mathbb{R}^{d\times d(M-1)}$ with the dimension of this subspace being less than $(M-1)d^{2}$, as it must satisfy a nontrivial linear equation. Consequently, the set of matrices $B_{2},...,B_{M}$ that satisfy Equation (B.27) has a Lebesgue measure zero in $[B_{2}:\cdots:B_{M}]\in\mathbb{R}^{d\times d(M-1)}$, implying that the set of matrices $B_{1},...,B_{M}$ satisfying this equation have a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$.
Suppose then $D_{t}P_{t}\neq I_{d}$ and denote the left side of Equation (B.26) as $C_{t}\in\mathbb{R}^{d\times d}$. By Condition (B) of Proposition 2, the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero. At each such time period $t$, the right side of Equation (B.26) is nonzero, thus, implying that also $C_{t}$ is not a matrix zeros. Equation (B.26) is thereby nontrivial and can be expressed as
@@
By the above discussion, for any single transformation $D_{t}P_{t}$ of $B_{y,t}$, the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.26) for any time period when all the entries of $(\alpha_{1,t},...,\alpha_{M,t})$ are nonzero has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Since there exists a finite number of such transformations and a countable union of measure zero sets has a measure zero, it follows that the set of matrices $B_{1},...,B_{M}$ that satisfy Equation (B.27) for all $t$ has a Lebesgue measure zero in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. This concludes that, almost everywhere in $[B_{1},...,B_{M}]\in\mathbb{R}^{d\times Md}$, the matrix $B_{y,t}(\boldsymbol{\tau},\boldsymbol{\iota})$ cannot be obtained by reordering or changing the signs of the columns of $B_{y,t}(\tilde{\boldsymbol{\tau}},\tilde{\boldsymbol{\iota}})$ whenever $\boldsymbol{\tau}_{m}\neq\tilde{\boldsymbol{\tau}}_{m}$ or $\boldsymbol{\iota}_{m}\neq\tilde{\boldsymbol{\iota}}_{m}$ for any $m=2,...,M$
Next, we make use of the above result together with Lemma 2 to show that $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times Md}$. Consider the impact matrix $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ with some invertible $(d\times d)$ matrices $B_{1},...,B_{M}$. By Lemma 2, $B_{y,t}$ is uniquely identified at each $t$ up to ordering and signs of its columns (conditionally on $\mathcal{F}_{t-1}$). Since the ordering and signs of the columns of $B_{1}$ are fixed (Condition (A) of Proposition 2), we then obtain from the above results that almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, reordering or changing the signs of the columns of $B_{2},...,B_{M}$ would, at some $t$, lead to an impact matrix that cannot obtained by reordering or changing the signs of the columns of $B_{y,t}$. Hence, models before and after reordering or changing the signs of the columns of any of $B_{2},...,B_{M}$ are not be observationally equivalent. The ordering and signs of the columns of $B_{1},...,B_{M}$ are thereby effectively fixed under Condition (A) of Proposition 2, implying that the ordering and signs of the columns of $B_{y,t}$ are also fixed for all $t$ (almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$). Thus, $B_{y,t}$ is uniquely identified at all $t$.
@@
Denote the $ij$th element of $\tilde{B}_{m}$ as $\tilde{b}_{ij}^{(m)}$. At each $t$, Equation (B.29) then defines a system of linear equations involving $d^{2}$ equations and $Md^{2}$ variables $\tilde{b}_{ij}^{(m)}$, $j,i=1,...,d$, $m=1,...,M$. The weights $\alpha_{m,t}$ are treated as constant coefficients here (some of which may be zero), as they are predetermined at each $t$ (by $\mathcal{F}_{t-1}$-measurability or nonrandomness), and the parameters in $B_{1},...,B_{M}$ are also constants here.
Each scalar equation in (B.29) is a linear equation in $M$ variables. Thus, each scalar equation in (B.29) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and $M$ variables. By Condition (3) of Lemma 2 (assumed in Proposition 2), there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent. Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td^{2}$ linear equations (B.29), $t=1,....,T$, in $Md^{2}$ variables. Since one solution is $\tilde{B}_{m}=B_{m}$ for all $m=1,...,M$, this is the unique solution, concluding that the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:\cdots:B_{M}]\in\mathbb{R}^{d\times dM}$, given that the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, are uniquely identified. For exogenous models, the identification of the transition weights is given. For logistic models, by Lemma 2, the transition weights weights are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$, concluding the proof.∎